# Inexact subgradient algorithm with a non-asymptotic convergence guarantee for copositive programming problems

### v1

@@ 5.2 Solving general copositive programming problems
[…]
$n$ $c$ $M$ GUnif SUnif GUnif-MILP GUnif SUnif
5 $5\text{\times}{10}^{-1}$ $2423$ $6.84\text{\times}{10}^{-01}$ $6.47\text{\times}{10}^{-01}$ $2.91\text{\times}{10}^{-03}$ $7.6\text{\times}{10}^{-03}$ $4.2\text{\times}{10}^{-03}$
$1\text{\times}{10}^{-1}$ $485$ $6.52\text{\times}{10}^{-01}$ $5.40\text{\times}{10}^{-01}$ $3.51\text{\times}{10}^{-02}$ $3.7\text{\times}{10}^{-03}$ $1.7\text{\times}{10}^{-03}$
$1\text{\times}{10}^{-2}$ $49$ $6.86\text{\times}{10}^{-01}$ $3.55\text{\times}{10}^{-01}$ $1.62\text{\times}{10}^{-03}$ $1.8\text{\times}{10}^{-03}$ $1.2\text{\times}{10}^{-03}$
$1\text{\times}{10}^{-3}$ 55 $2.89\text{\times}{10}^{-02}$ $2.52\text{\times}{10}^{-01}$ $7.16\text{\times}{10}^{-01}$ $1.5\text{\times}{10}^{-03}$ $1.4\text{\times}{10}^{-03}$
$\text{\times}{10}^{-4} 11$ $2.16\text{\times}{10}^{-02}$ $4.30\text{\times}{10}^{-01}$ $7.09\text{\times}{10}^{-01}$ $3.0\text{\times}{10}^{-03}$ $1.9\text{\times}{10}^{-03}$
10 $5\text{\times}{10}^{-1}$ 5 007 5035\,007\,503 $8.71\text{\times}{10}^{-01}$ $7.08\text{\times}{10}^{-01}$ $9.17\text{\times}{10}^{-02}$ $6.2\text{\times}{10}^{00}$ $4.9\text{\times}{10}^{-01}$
$1\text{\times}{10}^{-1}$ 1 001 5011\,001\,501 $8.69\text{\times}{10}^{-01}$ $7.14\text{\times}{10}^{-01}$ $9.43\text{\times}{10}^{-02}$ $3\text{\times}{10}^{00} 1.1$$1\text{\times}{10}^{-01}$
$1\text{\times}{10}^{-2}$ 100 151100\,151 $8.40\text{\times}{10}^{-01}$ $5.84\text{\times}{10}^{-01}$ $1.23\text{\times}{10}^{-01}$ $1.6\text{\times}{10}^{-01}$ $2.1\text{\times}{10}^{-02}$
$1\text{\times}{10}^{-3}$ 10 01610\,016 $6.80\text{\times}{10}^{-01}$ $5.80\text{\times}{10}^{-01}$ $2.83\text{\times}{10}^{-01}$ $3.4\text{\times}{10}^{-02}$ $4.8\text{\times}{10}^{-03}$
$1\text{\times}{10}^{-4}$ $1002$ $6.18\text{\times}{10}^{-01}$ $4.03\text{\times}{10}^{-01}$ $3.45\text{\times}{10}^{-01}$ $5.4\text{\times}{10}^{-03}$ $1.7\text{\times}{10}^{-03}$
[…]
Following Section 5.1, we used MILP and Grid to solve a standard quadratic programming problem that appears as a subproblem at each iteration of Algorithm 1. We set $\alpha=0$ when using MILP to solve subproblems and $\alpha=1$ when using Grid. For such $\alpha$, we set $\epsilon$ so that $\epsilon(1+\alpha)$ was $2$ or $0.2$. When using Grid to solve a standard quadratic programming problem of the form (4.7) at each iteration, we set $r$ to the minimum positive integer satisfying (4.12).
@@
2 5 MILP $6.4\text{\times}{10}^{00}$ $3.02\text{\times}{10}^{00}$ $6.3\text{\times}{10}^{-02}$
@@
0.2 5 MILP $6.1\text{\times}{10}^{02}$ $5.77\text{\times}{10}^{-01}$ $9.8\text{\times}{10}^{-01}$
Grid $2.4\text{\times}{10}^{03}$ $2.88\text{\times}{10}^{-01}$ $4.6\text{\times}{10}^{-01}$
10 MILP $1.0\text{\times}{10}^{03}$ $6.64\text{\times}{10}^{-01}$ $2.2\text{\times}{10}^{00}$
Grid >>3600 >>3600 >>3600
50 MILP $1.5\text{\times}{10}^{03}$ $3.81\text{\times}{10}^{-01}$ $3.8\text{\times}{10}^{01}$
Grid >>3600 >>3600 >>3600
100 MILP $1.8\text{\times}{10}^{03}$ $4.06\text{\times}{10}^{-01}$ $3.1\text{\times}{10}^{02}$
Grid >>3600 >>3600 >>3600
500 GUnif $1.1\text{\times}{10}^{04}$ $7.89\text{\times}{10}^{01}$ $1.6\text{\times}{10}^{03}$
@@
The number of samples in GUnif is not consistent with Proposition 4.5 and is much smaller than the number of points in the grid.*2*2*2If we set $r$ to the minimum positive integer satisfying (4.24), the number of points in $\Delta_{r}^{n-1}$ sometimes exceeded realmax in MATLAB, which is the largest finite floating-point number in IEEE double precision and is approximately $1.7977\text{\times}{10}^{308}$. The average objective values obtained using GUnif were rather coarse. However, when MILP and Grid take too long to solve problems, employing GUnif to obtain approximate solutions may be a viable alternative.
@@ 5.3 Testing for non-complete positivity of exceptional doubly nonnegative matrices
A real symmetric matrix is said to be doubly nonnegative if it is positive semidefinite and entrywise nonnegative. It is well known that every completely positive matrix is doubly nonnegative, and the converse holds if and only if the order is less than or equal to $4$; see [25, page 101] and [38]. We can easily check whether a given matrix is doubly nonnegative, so we are interested in deciding whether a given doubly nonnegative matrix is completely positive. In Theorem 4.8, we provided a sufficient condition for a given matrix not to be completely positive. In this subsection, we demonstrate that this theorem is effective for doubly nonnegative matrices that are not completely positive. For convenience, following the terminology of [53], we call a doubly nonnegative matrix that is not completely positive exceptional.*3*3*3Burer, Anstreicher, and Dür [14] refer to an exceptional doubly nonnegative matrix as a bad matrix.
@@
Table 6: The results of detecting non-complete positivity for the exceptional doubly nonnegative matrices presented in [4, Appendix B]. In the “Non-CP” column, “Y” denotes successful detection, whereas “N” indicates failure to detect non-complete positivity.
@@
First, we tried to detect non-complete positivity for the ten $6\times 6$ exceptional doubly nonnegative matrices presented in [4, Appendix B]. We normalized each matrix and used the normalized matrix as the input matrix $\bm{C}$ in (4.25). In this experiment, we set the constant $t$ that appears in (4.33) to $55$. Table 6 shows the results of detecting non-complete positivity for the ten matrices. Except for the matrix extremal_rand_5, we successfully detected non-complete positivity for all the other matrices.
@@
where $c_{i}=1$ if $i=1$ and $c_{i}=\sqrt{2}$ if $i\geq 2$. Note that we can calculate (5.1) analytically; see [53, Equation (2.3)]. Then, inspired by the method for constructing exceptional doubly nonnegative matrices presented in [53, Equation (1.12)], we solved the following semidefinite feasibility problem:
@@
However, we note that the choice of the constant $t$ that appears in (4.33) is crucial. The smaller $t$ is, the fewer the iterations, but at the same time the harder it is to detect non-complete positivity for a matrix. In fact, if we set $t$ to $5$, we could not detect non-complete positivity for any of the exceptional doubly nonnegative matrices we used in this subsection. We need to consider how to set $t$ to balance the number of iterations and the performance of detecting non-complete positivity.
@@ 6 Conclusion
In this paper, based on the subgradient algorithm in [43, Equation (3.2.24)], which is applicable to convex programming problems with a single nonsmooth functional constraint, we proposed an algorithm with a non-asymptotic convergence guarantee to solve copositive programming problems. In contrast to the algorithm in [43, Equation (3.2.24)], the proposed algorithm allows us to solve the subproblem, which is a standard quadratic programming problem, inexactly at each iteration. We discussed how to solve the standard quadratic programming problem exactly and inexactly. Through the numerical experiments, we saw that for small-scale problems, the exact method for standard quadratic programming via mixed-integer linear programming is acceptable. For medium-scale problems, the exact method may be effective in some cases, whereas the inexact deterministic method obtained by discretizing the unit simplex using a regular grid may be effective in others. Note, however, that the inexact method has the advantage that it does not require external solvers. For large-scale problems where the above two methods take too long to solve, uniform sampling from the regular grid might be another option to solve standard quadratic programming problems inexactly, although its solution accuracy is coarse.

### v2

@@ 5.2 Solving general copositive programming problems
Following Section 5.1, we used MILP and Grid to solve a standard quadratic programming problem that appears as a subproblem at each iteration of Algorithm 1. When using MILP to solve subproblems, we set $\alpha=0$ and $\epsilon=2$, and when using Grid, we set $\alpha=1$ and $\epsilon=1$, so that $\epsilon(1+\alpha)=2$ in both cases. When using Grid to solve a standard quadratic programming problem of the form (4.7) at each iteration, we set $r$ to the minimum positive integer satisfying (4.12).
@@
5 MILP $6.4\text{\times}{10}^{00}$ $3.02\text{\times}{10}^{00}$ $6.3\text{\times}{10}^{-02}$
@@
The number of samples in GUnif is not consistent with Proposition 4.5 and is much smaller than the number of points in the grid.*2*2*2If we set $r$ to the minimum positive integer satisfying (4.25), the number of points in $\Delta_{r}^{n-1}$ sometimes exceeded realmax in MATLAB. The average objective values obtained using GUnif were rather coarse. However, when MILP and Grid take too long to solve problems, employing GUnif to obtain approximate solutions may be a viable alternative.
@@ 5.3 Testing for non-complete positivity of exceptional doubly nonnegative matrices
A real symmetric matrix is said to be doubly nonnegative if it is positive semidefinite and entrywise nonnegative. It is well known that every completely positive matrix is doubly nonnegative, and the converse holds if and only if the order is less than or equal to $4$; see [27, page 101] and [39]. We can easily check whether a given matrix is doubly nonnegative, so we are interested in deciding whether a given doubly nonnegative matrix is completely positive. In Theorem 4.8, we provided a sufficient condition for a given matrix not to be completely positive. In this subsection, we demonstrate that this theorem is effective for doubly nonnegative matrices that are not completely positive. For convenience, following the terminology of [54], we call a doubly nonnegative matrix that is not completely positive exceptional.*3*3*3Burer, Anstreicher, and Dür [14] refer to an exceptional doubly nonnegative matrix as a bad matrix.
@@
Table 5: The results of detecting non-complete positivity for the exceptional doubly nonnegative matrices presented in [5, Appendix B]. In the “Non-CP” column, “Y” denotes successful detection, whereas “N” indicates failure to detect non-complete positivity.
@@
First, we tried to detect non-complete positivity for the ten $6\times 6$ exceptional doubly nonnegative matrices presented in [5, Appendix B]. We normalized each matrix and used the normalized matrix as the input matrix $\bm{C}$ in (4.26). In this experiment, we set the constant $t$ that appears in (4.34) to $55$. Table 5 shows the results of detecting non-complete positivity for the ten matrices. Except for the matrix extremal_rand_5, we successfully detected non-complete positivity for all the other matrices.
@@
where $c_{i}=1$ if $i=1$ and $c_{i}=\sqrt{2}$ if $i\geq 2$. Note that we can calculate (5.1) analytically; see [54, Equation (2.3)]. Then, inspired by the method for constructing exceptional doubly nonnegative matrices presented in [54, Equation (1.12)], we solved the following semidefinite feasibility problem:
@@
However, the choice of the constant $t$ in (4.34) is crucial, and in the two experiments we selected $t$ by trial and error. The smaller $t$ is, the fewer the iterations, but at the same time the harder it is to detect non-complete positivity for a matrix. In the first experiment, setting $t=50$ failed to detect non-complete positivity for the matrix extremal_rand_1. In the second experiment, setting $t=5$ failed to detect non-complete positivity for any of the exceptional doubly nonnegative matrices. We need to consider how to set $t$ to balance the number of iterations and the performance of detecting non-complete positivity.
@@ 6 Conclusion
In this paper, based on the subgradient algorithm in [43, Equation (3.2.24)], which is applicable to convex programming problems with a single nonsmooth functional constraint, we proposed an algorithm with a non-asymptotic convergence guarantee to solve copositive programming problems. In contrast to the algorithm in [43, Equation (3.2.24)], the proposed algorithm allows us to solve the subproblem, which is a standard quadratic programming problem, inexactly at each iteration. We discussed how to solve the standard quadratic programming problem exactly and inexactly.
Through the numerical experiments, we identified trends indicating which method is preferable under which circumstances for solving standard quadratic programming problems. For small-scale problems, the exact method for standard quadratic programming via mixed-integer linear programming is acceptable. For medium-scale problems, the exact method may be effective in some cases, whereas the inexact deterministic method obtained by discretizing the standard simplex using a regular grid may be effective in others. Note, however, that the inexact method has the advantage that it does not require external solvers. For large-scale problems where the above two methods take too long to solve, uniform sampling from the regular grid might be another option to solve standard quadratic programming problems inexactly, although its solution accuracy is coarse.